OpenStudy (anonymous):

What's the real solution for the equation a^x+1= b/a? (consider a>0, a different from 1 and b>0) a) loga (b) b) loga (b) - 2 c) loga (b) + 2 d) loga (b+1) Can someone tell me only how to start solving this?

OpenStudy (anonymous):

i take it this is \[a^{x+1}=\frac{b}{a}\] yes?

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

if so then you can write \[a\times a^{x+1}=b\] \[a^{x+2}=b\] \[x+2=\log_a(b)\] \[x=\log_a(b)-2\]

OpenStudy (anonymous):

a (a^x+1) shouldnt be a^2x+2 ?

OpenStudy (anonymous):

@myininaya could you explain that to me pls?

myininaya (myininaya):

?

OpenStudy (anonymous):

I didn't understand the answer. how come a (a^x+1) shouldnt be a^2x+2 ?

myininaya (myininaya):

We want to show \[a(a^{x+1})=a^{2x+2}?\]

OpenStudy (anonymous):

Yes

myininaya (myininaya):

\[a \cdot a^{x+1}=a^{1} \cdot a^{x+1}=a^{1+(x+1)}=a^{x+1+1}=a^{x+2}\]

myininaya (myininaya):

Same base and we are multiply add exponents!

OpenStudy (anonymous):

is it wrong to do a(a^x+1) => a²^(x+1) and then a^2x+2 ?

myininaya (myininaya):

I don't understand those steps and yes that is wrong

OpenStudy (anonymous):

Thank you so much myini!!